A rectangle is 3cm longer than it is wide. The diagonal is 15cm. Find the dimensions of the rectangle.
Answer by checkley77(7072)

L=W+3
L^2+W^2=15^2
(W+3)^2+W^2=225
W^2+6W+9+W^2=225
2W^2+6W+9-225=0
2W^2+6W-216=0
2(W^W+3W-108)=0
2(W+12)(W-9)=0
W-9=0
W=9 ANS. FOR THE WIDTH.
L=9+3=12 ANS. FOR THE LENGTH.
pROOF:
9^2+12^2=1562
81+144=225
225=225